OpenStudy (anonymous):

Simplify & Solve: x+x+10+2x-10=180

OpenStudy (anonymous):

x+x+10+2x-10=180 4x+10=180 4x=170 x=42.5

OpenStudy (anonymous):

x+x+10+2x-10=180 4x = 180 x = 45

OpenStudy (anonymous):

woops 3pwood is right I forgot about the positive ten x+x+10+2x-10=180 4x=180 x=45